# LLM as OS (llmao), Agents as Apps: Envisioning AIOS, Agents and the AIOS-Agent Ecosystem

### 06 Dec 2023 v1

@@ 3.4.2 Tool-Driver and Tool-API
Utilizing the inherent zero-shot and few-shot learning capabilities of LLMs (Radford et al., 2019; Brown et al., 2020), agents can gain insights about tools through zero-shot prompts that elucidate tool functionalities and parameters, or few-shot prompts offering demonstrations of particular tool usage scenarios and methodologies (Schick et al., 2023; Parisi et al., 2022). Usually, a single tool is inadequate for complex tasks. Hence, agents must adeptly decompose these tasks into manageable subtasks, where their understanding of the tools is pivotal. After understanding individual tools, LLMs should determine their application in addressing complex tasks. One approach is to generate actions by extracting pertinent information from memory relevant to the current task. For instance, Generative Agents (Park et al., 2023) maintain a memory stream, consulting it for recent, pertinent, and crucial information before each action to guide their decisions. In GITM (Zhu et al., 2023), to achieve specific sub-goals, the agent probes its memory to identify if any similar tasks have been successfully executed before. If so, it replicates those successful actions for the current task. In collaborative frameworks such as ChatDev (Qian et al., 2023a) and MetaGPT (Hong et al., 2023), agents engage in mutual communication, retaining the conversation history in their memories. Another strategy involves executing a pre-formulated plan. For example, in DEPS (Wang et al., 2023a), given a task, if there are no indicators of the plan’s failure, the agent proceeds with actions based on that plan.
@@ 4.1 Agents as Applications
On top of the LLMAO-based AIOS, a diverse scope of AI Agent applications can be developed, akin to traditional OS-based applications such as browsers and photo-editing softwares. These Agent Applications (AAPs) generally comprise three components: the agent profile, an accessible external database, and task-specific tools. On one hand, the agent profiles are written into the prompt and used to indicate the agent functionality, influencing the LLM behaviors to exhibit certain roles such as a coder agent, a teacher agent, or a travel planning agent, as described in sources such as Qian et al. (2023a); Li et al. (2023); Ge et al. (2023). Typical agent profiles may encompass basic information such as age, gender, and career (Park et al., 2023), or psychology information reflecting the personalities of the agents (Serapio-García et al., 2023), as well as social information detailing the relationships between agents (Li et al., 2023). The nature of the agent’s profile is tailored to the application’s context; for example, applications focused on human cognitive processes will emphasize psychological information. Furthermore, the inclusion of an external database and toolset within the prompt enhances the LLM’s interaction with the external world, specifying elements such as file locations and tool descriptions.
As illustrated in Figure 4, by merging the LLMAO layer with the OS layer, Hardware layer, and the Agent Application layer, we can establish an autonomous LLMAO-based Agent system. This system will respond to natural language instructions from users, capable of executing a variety of tasks through its interactions with the environment and its inherent knowledge.
@@ 4.2 Natural Language Programming for Agents
Natural Language Programming (NLProg) acts as a crucial intermediary between LLMAO and Agents within the AIOS-Agent ecosystem. This innovative approach allows average users to easily program Agent Applications (AAPs) using natural language. This development democratizes the creation and accessibility of computer software, representing a significant shift from the traditional OS-APP ecosystem. Traditionally, desktop or mobile applications (APPs) required programming by skilled software developers using professional programming languages. In contrast, NLProg empowers even those without formal training in these languages to develop applications.
@@ 4.3 The Ecosystem
5 LLMAO in Practice: AI Agents
In this section, we present a comprehensive overview of the current applications of LLMAO-based agents, aiming to provide a wide-ranging perspective on their practical deployment scenarios. Specifically, we mainly introduce three scenarios: single agent applications, multi-agent applications, and human-agent applications.
@@ 6.1.1 Memory Management
LLMs have revolutionized abilities, but are constrained by limited context windows, hindering their utility in tasks such as extended conversations and document analysis. To mitigate the limited context window problem in LLMs, approaches inspired by the above operating system memory management methods can be developed. A swapping mechanism, similar to OS swapping to external storage, can be implemented in LLMs to temporarily store inactive parts of their context, effectively enlarging the context window. This would require efficient retrieval systems to minimize latency. For example, MemGPT (Packer et al., 2023) intelligently manages different memory tiers in order to effectively provide extended context within the LLM’s limited context window, and utilizes interrupts to manage control flow between itself and the user. Additionally, mirroring OS memory sharing, LLMAO can share context or learned patterns across different LLMAO-based Agent instances, reducing redundancy and allowing access to a larger shared knowledge pool for agents. Finally, drawing from memory disaggregation in modern OS, agents can leverage networked memory resources, enabling them to access and process data stored across multiple LLMAOs, thus expanding their abilities significantly. Each of these strategies, while offering potential solutions, also presents unique challenges such as managing latency, ensuring consistency in shared contexts, and handling the complexities of distributed memory systems.
@@ 6.2 Communication
Overall, the use of a large language model as an interpreter (LLM as Interpreter), which can translate natural language described plans to DSL described plans, can be an important direction to create task-solving plans, and to bridge LLM as OS (LLMAO) and Agent Applications (AAP). The development of a DSL for such plans could further streamline and standardize the process, though it would come with its own set of challenges. This approach has the potential to make complex task execution more accessible and efficient, paving the way for more advanced AI systems in the future.
@@ 6.3 Security
Similarly, adversarial attacks have been the subject of extensive study in LLM research (Wei et al., 2023; Zou et al., 2023; Qi et al., 2023a; Yang et al., 2023c), representing a significant threat to the security of AIOS. Furthermore, research (Yang et al., 2023b; Qi et al., 2023b) has revealed that an aligned LLM can be broken using a very small dataset comprising only a few hundred data points. This vulnerability is not only a significant concern in terms of system integrity but also raises alarming implications for the safety of agents interacting with these systems. How to be robust and defend against such attacks has yet to be studied in the context of LLMAO, AIOS, and Agents.
Looking into the future, as agent applications in AIOS evolve to be programmed by natural language, the intricacy of scanning the code of such applications will increase due to the expansive and less structured nature of natural language compared to the constrained syntax of programming languages. This necessitates a robust and effective scanning tool for AIOS agents. Moreover, fuzzing can be seen as an initial step towards creating a Red-teaming dataset for LLMAO-based Agents. For instance, Ruan et al. (2023b) have developed a multi-agent system that generates red-teaming scenarios for LLM-based agents, using GPT-4 to simulate adversarial environments within textual scenarios. The study provides a rich array of scenarios that serve as a valuable resource for future research into the alignment of LLM-based Agents.

### 09 Dec 2023 v2

@@ 3.4.2 Tool-Driver and Tool-API
Utilizing the inherent zero-shot and few-shot learning capabilities of LLMs (Radford et al., 2019; Brown et al., 2020), agents can gain insights about tools through zero-shot prompts that elucidate tool functionalities and parameters, or few-shot prompts offering demonstrations of particular tool usage scenarios and methodologies (Schick et al., 2023; Parisi et al., 2022). Usually, a single tool is inadequate for complex tasks. Hence, agents must adeptly decompose these tasks into manageable subtasks, where their understanding of the tools is pivotal. After understanding individual tools, LLMs should determine their application in addressing complex tasks. One approach is to generate actions by extracting pertinent information from memory relevant to the current task. For instance, Generative Agents (Park et al., 2023) maintain a memory stream, consulting it for recent, pertinent, and crucial information before each action to guide their decisions. In GITM (Ghost in the Minecraft) (Zhu et al., 2023), to achieve specific sub-goals, the agent probes its memory to identify if any similar tasks have been successfully executed before. If so, it replicates those successful actions for the current task. In collaborative frameworks such as ChatDev (Qian et al., 2023a) and MetaGPT (Hong et al., 2023), agents engage in mutual communication, retaining the conversation history in their memories. Another strategy involves executing a pre-formulated plan. For example, in DEPS (Describe, Explain, Plan and Select) (Wang et al., 2023a), given a task, if there are no indicators of the plan’s failure, the agent proceeds with actions based on that plan.
@@ 4.1 Agents as Applications
On top of the LLMOS-based AIOS, a diverse scope of AI Agent applications can be developed, akin to traditional OS-based applications such as browsers and photo-editing softwares. These Agent Applications (AAPs) generally comprise three components: the agent profile, an accessible external database, and task-specific tools. On one hand, the agent profiles are written into the prompt and used to indicate the agent functionality, influencing the LLM behaviors to exhibit certain roles such as a coder agent, a teacher agent, or a travel planning agent, as described in sources such as Qian et al. (2023a); Li et al. (2023); Ge et al. (2023). Typical agent profiles may encompass basic information such as age, gender, and career (Park et al., 2023), or psychology information reflecting the personalities of the agents (Serapio-García et al., 2023), as well as social information detailing the relationships between agents (Li et al., 2023). The nature of the agent’s profile is tailored to the application’s context; for example, applications focused on human cognitive processes will emphasize psychological information. Furthermore, the inclusion of an external database and toolset within the prompt enhances the LLM’s interaction with the external world, specifying elements such as file locations and tool descriptions.
As illustrated in Figure 4, by merging the LLMOS layer with the OS layer, Hardware layer, and the Agent Application layer, we can establish an autonomous LLMOS-based Agent system. This system will respond to natural language instructions from users, capable of executing a variety of tasks through its interactions with the environment and its inherent knowledge.
Figure 4: An illustration of LLMOS-based AI Agent.
@@ 4.2 Natural Language Programming for Agents
Natural Language Programming (NLProg) acts as a crucial intermediary between LLMOS and Agents within the AIOS-Agent ecosystem. This innovative approach allows average users to easily program Agent Applications (AAPs) using natural language. This development democratizes the creation and accessibility of computer software, representing a significant shift from the traditional OS-APP ecosystem. Traditionally, desktop or mobile applications (APPs) required programming by skilled software developers using professional programming languages. In contrast, NLProg empowers even those without formal training in these languages to develop applications.
@@ 4.3 The Ecosystem
5 LLMOS in Practice: AI Agents
In this section, we present a comprehensive overview of the current applications of LLMOS-based agents, aiming to provide a wide-ranging perspective on their practical deployment scenarios. Specifically, we mainly introduce three scenarios: single agent applications, multi-agent applications, and human-agent applications.
@@ 6.1.1 Memory Management
LLMs have revolutionized abilities, but are constrained by limited context windows, hindering their utility in tasks such as extended conversations and document analysis. To mitigate the limited context window problem in LLMs, approaches inspired by the above operating system memory management methods can be developed. A swapping mechanism, similar to OS swapping to external storage, can be implemented in LLMs to temporarily store inactive parts of their context, effectively enlarging the context window. This would require efficient retrieval systems to minimize latency. For example, MemGPT (Packer et al., 2023) intelligently manages different memory tiers in order to effectively provide extended context within the LLM’s limited context window, and utilizes interrupts to manage control flow between itself and the user. Additionally, mirroring OS memory sharing, LLMOS can share context or learned patterns across different LLMOS-based Agent instances, reducing redundancy and allowing access to a larger shared knowledge pool for agents. Finally, drawing from memory disaggregation in modern OS, agents can leverage networked memory resources, enabling them to access and process data stored across multiple LLMOSes, thus expanding their abilities significantly. Each of these strategies, while offering potential solutions, also presents unique challenges such as managing latency, ensuring consistency in shared contexts, and handling the complexities of distributed memory systems.
@@ 6.2 Communication
Overall, the use of a large language model as an interpreter (LLM as Interpreter), which can translate natural language described plans to DSL described plans, can be an important direction to create task-solving plans, and to bridge LLM as OS (LLMOS) and Agent Applications (AAP). The development of a DSL for such plans could further streamline and standardize the process, though it would come with its own set of challenges. This approach has the potential to make complex task execution more accessible and efficient, paving the way for more advanced AI systems in the future.
@@ 6.3 Security
Similarly, adversarial attacks have been the subject of extensive study in LLM research (Wei et al., 2023; Zou et al., 2023; Qi et al., 2023a; Yang et al., 2023c), representing a significant threat to the security of AIOS. Furthermore, research (Yang et al., 2023b; Qi et al., 2023b) has revealed that an aligned LLM can be broken using a very small dataset comprising only a few hundred data points. This vulnerability is not only a significant concern in terms of system integrity but also raises alarming implications for the safety of agents interacting with these systems. How to be robust and defend against such attacks has yet to be studied in the context of AIOS, LLMOS, and Agents.
Looking into the future, as agent applications in AIOS evolve to be programmed by natural language, the intricacy of scanning the code of such applications will increase due to the expansive and less structured nature of natural language compared to the constrained syntax of programming languages. This necessitates a robust and effective scanning tool for AIOS agents. Moreover, fuzzing can be seen as an initial step towards creating a Red-teaming dataset for LLMOS-based Agents. For instance, Ruan et al. (2023b) have developed a multi-agent system that generates red-teaming scenarios for LLM-based agents, using GPT-4 to simulate adversarial environments within textual scenarios. The study provides a rich array of scenarios that serve as a valuable resource for future research into the alignment of LLM-based Agents.